# Durham, UK

### v1

@@ Abstract
A consistent treatment of the quantum field theory of an axion-like particle (ALP) interacting with Standard Model fields requires to account for renormalisation group running and matching to the low-energy theory. Quantum sensor experiments designed to search for very light ALPs are particularly sensitive to these effects because they probe large values of the decay constant for which running effects become important. In addition, while linear axion interactions are set by its pseudoscalar nature, quadratic interactions are indistinguishable from quadratic interactions of scalars. We show how the Wilson coefficients of linear and quadratic axion interactions are related, including running effects above and below the QCD scale and provide a comprehensive analysis of the sensitivity of current and future experiments. We identify the reach of different experiments for the case of ALP dark matter and comment on how it could be distinguished from the case where it is not the dark matter. We present novel search strategies to observe quadratic ALP interactions via fifth force searches, haloscopes, helioscopes and quantum sensors. We emphasize the nonlinear behaviour of the ALP field close to the surface of the earth and point out which experimental results can be trusted in a regime where the ALP background field has unphysical values.
@@ 1 Introduction
Quantum sensors have significantly enhanced their sensitivity over the past decade, creating new opportunities to explore fundamental physics questions. These advancements are particularly valuable in testing the effects of light new physics through more precise quantum sensor experiments Safronova:2017xyt ; Antypas:2022asj . A prime example of this light new physics is pseudoscalars, which emerge in theories where an approximate global symmetry is spontaneously broken Peccei:1977hh ; Peccei:1977ur ; Weinberg:1977ma ; Wilczek:1977pj . These are referred to as pseudo-Nambu Goldstone bosons, or more commonly, axion-like particles (ALP). The goal of this paper is to translate already performed and anticipated future precision measurements with quantum sensors into a sensitivity range for ALPs. We take the full effective ALP Lagrangian into account, include higher order effects when relevant and perform a consistent calculation of the axion couplings at different scales accounting for renormalisation effects. We stress the importance of quadratic ALP couplings that are formally subleading, but induce spin-independent effects that are strongly constrained from experiments that have not been considered for pseudoscalar interactions.
@@ 2.2 Quadratic ALP Interactions
where we introduced $\hat{m}=(m_{u}+m_{d})/2$ and $\Delta_{m}=(m_{u}-m_{d})/2$ to make it easier to identify the isospin-breaking terms and $\tau_{a}=m_{a}^{2}/m_{\pi}^{2}$. The sign of $\delta_{\pi}(a)$ plays a crucial role in determining the environmental effects of a massive body influencing the ALP field value. Our result (13) agrees with Kim:2022ype ; Kim:2023pvt in the limit $m_{a}\to 0$, but disagrees with the result used in Beadle:2023flm .
@@
$\alpha^{\text{eff}}(a)$ $=\Big{(}1+\delta_{\alpha}(a)\Big{)}\alpha\,\qquad\text{with}\qquad\delta_{\alpha}(a)=\frac{1}{12\pi}\left(1-32c_{1}\frac{m_{\pi}^{2}}{M_{N}}\right)\,\delta_{\pi}(a)$ (24)
In principle, there is also a tree-level contribution from the first diagram on the right-hand side of Figure 1, which is suppressed by another factor of $\alpha$ with respect to the contributions in (24). Shift symmetry breaking in quadratic ALP-lepton couplings is a loop effect because the tree-level diagram in the lower row of Figure 1 results in a momentum suppressed contribution to a higher order operator with additional derivatives $(\partial a)^{2}$. The leading contribution to the quadratic ALP electron coupling is a result of photon loops Kim:2023pvt
@@ 2.5 Quadratic ALP couplings and the ALP background field
$\left(\partial_{t}^{2}-\Delta+\bar{m}_{a}^{2}(r)\right)a=J_{\text{source}}(r)$ (45)
where $J_{\text{source}}(r)$ is suppressed by CP violating parameters or proportional to $\gamma_{5}$, and the effective mass term can be written as
@@
$a(r,t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{a}}\cos\big{(}m_{a}t\big{)}\bigg{[}1-Z(\delta_{i})J_{\pm}\left(\sqrt{3|Z(\delta_{i})|}\right)\frac{R_{\text{source}}}{r}\bigg{]}$ (47)
and
@@
This can be inferred from the form of the pion mass shift $\delta_{\pi}<0$ as given in (13), which is strictly negative for any sign of $c_{GG}$. In the presence of ALP-quark couplings, this is not guaranteed, but any such contribution is heavily suppressed by the small ALP mass. As a consequence, it follows from (18), (22), (24) and (26) that for all coefficients $\delta_{i}<0$, $i=\pi,M_{N},\Delta M,\alpha,e$. At the same time, the charges in (2.3) are positive in the case of an ALP interacting with gluons, for which the sole negative contributions are heavily suppressed since the ‘dilatonic’ charges $Q_{\hat{m}},Q_{\Delta M}\gg Q_{\alpha},Q_{e}$. Note that this is different for a scalar with quadratic couplings or even a more general model in which the sign of (51) can be more model-dependent. Note also, that (47) depends on the boundary conditions of the ALP field Banerjee:2022sqg .
@@
As a result, the ALP field (47) diverges close to the surface of the earth with crucial consequences for measurements of varying fundamental constants that are sensitive to the ALP field value squared, as well as for experiments measuring nuclear magnetic resonances, electric dipole oscillations Hill:2015kva ; Chu:2018avg and atomic spin precession in the ALP background field Graham:2017ivz ; Abel:2017rtm ; Agrawal:2022wjm ; Dror:2022xpi . In contrast to quadratically interacting scalars, fundamental constants cannot change sign in this parameter space Stadnik:2021qyf ; Banerjee:2022sqg .
@@
Figure 2: Bounds from fifth forces on ALP-gluon coupling at the UV scale.
@@ 3.1 Fifth forces
An important effect that can distinguish between very light ALP dark matter and an ALP field that does not make up a significant amount of dark matter is the acceleration experienced by test masses close to massive bodies like planets, e.g. tests of violations of the equivalence principle. Forces caused by axion fields that are not dark matter are captured by the potentials (55) and (58). The acceleration of a test mass $m_{T}(a)$ in the ALP background field $a(x)$ follows from the equations of motion from the low energy Lagrangian
@@
Tests of the equivalence principle are particularly sensitive to non-universal ALP couplings. We stress again that parts of this parameter space is excluded due to the nonlinearity of the ALP field close to earth, which is not a concern for fifth force searches in the case that ALPs are not the dark matter. Future tests with muonic atoms could significantly improve the sensitivity of fifth-force searches Stadnik:2022gaa .
In Figure 2 we compare constraints from different searches for fifth forces induced by ALP exchange assuming a model with one coefficient, $c_{GG}$, in the UV. Searches for so-called Casimir-less forces are shown in blue Klimchitskaya:2015zpa ; Bauer:2023czj . The dominant contribution to this force is induced by the spin-independent potential (58). Torsion balances probe the same spin-independent force and the constraint from Adelberger:2006dh is shown in orange. Spin-dependent (dipole-dipole) interactions have been probed with comagnetometers using spin-polarised K and 3He atoms Vasilakis:2008yn and 87Rb Almasi:2018cob . The purple region labeled Ramsay scattering is excluded from measurements of spin-dependent forces in molecular H2 Ramsey:1979bzw . Future measurements of spin-dependent interactions exploiting nuclear magnetic resonances (NMR) can significantly improve the current sensitivity as shown in green dashed line in Figure 2 Arvanitaki:2014dfa ; Aybas:2021nvn . Note, that nuclear magnetic resonances are sensitive to CP-violating ALP couplings. In the case the ALP constitutes dark matter, the strongest constraint from fifth force searches are obtained by searches for violations of the equivalence principle. The MICROSCOPE collaboration measured the force necessary to keep two test masses in relative equilibrium on a satellite orbiting earth  Touboul:2017grn ; Touboul:2022yrw . The corresponding bound is shown in gray in Figure 2, but we emphasize that this parameter space is affected by the uncontrolled behaviour of the axion field value discussed in Section 2.5.
@@ 4.1 Quantum Clocks
where $\delta_{p}(a)$ is defined similar to $\delta_{N}(a)$ in (18) and denotes shift in proton mass. Plugging (13), (18), (24) and (26) into (75), one can express $\frac{\delta\nu_{A/B}}{\nu_{A/B}}$in terms of the low-energy Lagrangian parameters and the ratio $a^{2}/f^{2}$. For an oscillating ALP dark matter field, the quadratic ALP field evolves as
@@ 4.1.1 Current limits: microwave and optical clocks
Limits are obtained from the Rb/Cs atomic fountain clock, which compares the transition frequencies between different hyperfine levels in the two ground state atoms 87Rb and 133Cs. The frequency ratio measurements are sensitive to the variations in all three fundamental quantities in (75Hees:2016gop ; Flambaum:2023bnw . The measurements are also sensitive to very low frequencies corresponding to ALP mass $\sim 10^{-20}$ eV and below, due to the long time-span of the experiment. Throughout the frequency range of the experiment, the coherence time of the oscillation, $\tau_{a}=2\pi/\omega/(v/c)^{2}$ ($v/c\sim 10^{-3}$, $v$ is the virial velocity of dark matter in the galaxy) remains larger than the total timescale of measurement $T$. Although $T<<\tau_{c}$ maintains a coherent signal throughout the experiment, it also induces stochastic fluctuations in the dark matter amplitude Centers:2019dyn . Therefore, it is no longer accurate to assume the oscillation amplitude as simply $\sqrt{2\rho_{\rm DM}}/m_{a}$ because this will rather be a random variable with a sampling probability as determined by the distribution of the stochastic ALP field. This leads to corrections on the dark matter coupling that can vary from a factor of 2 up to 10.
We show the bounds from Rb/Cs in Figure 6 in the ${c_{GG}/f}$ ( UV scale) vs. $m_{a}$ plane where the coverage is for $10^{-24}\,\text{-}\,10^{-18}\ {\rm eV}$ ALP mass [^5] For dark matter below $\sim 10^{-22}$ eV, cosmological constraints prevent the assumption that dark matter density of the Universe is entirely due to the ALP field because the de-Broglie wavelength of the ALP exceeds the size of a dwarf galaxy 2168507 . Therefore the low-frequency limits should be rescaled because $\rho_{\rm DM}$ should be a fraction of the total dark matter density. We show the limits assuming ALP constitutes 100% of dark matter density.. The constant shift in the quadratic variation in the axion field is ignored Kim:2022ype . The sensitivity coefficients $k_{\alpha},\,k_{e},\,k_{q}$ and the frequency comparison data are obtained from Ref. Hees:2016gop . The local dark matter density is assumed to be ${\rm\rho_{DM}=0.4\ GeV/cm^{3}}$ [^6] The bounds are shown without including the stochastic fluctuation factor appearing in dark matter amplitude as the exact correction factor is unclear for a pseudoscalar dark matter..
@@ 4.3 Optical/laser Interferometers
$\frac{\delta n(t)}{n}$ $\approx 10^{-2}\bigg{(}\frac{\delta\omega(t)}{\omega_{0}}-$
$\quad[2\delta\alpha+\delta_{e}+(\delta_{e}-\delta_{p})/2-0.024(\delta_{q}-\delta_{p})]\,\frac{\rho_{\text{DM}}}{m_{a}^{2}f^{2}}\cos\,(2\omega_{a}t)\bigg{)}$ (82)

### v2

@@ Abstract
A consistent treatment of the quantum field theory of an axion-like particle (ALP) interacting with Standard Model fields requires to account for renormalisation group running and matching to the low-energy theory. Quantum sensor experiments designed to search for very light ALPs are particularly sensitive to these effects because they probe large values of the decay constant for which running effects become important. In addition, while linear axion interactions are set by its pseudoscalar nature, quadratic interactions are indistinguishable from scalar interactions. We show how the Wilson coefficients of linear and quadratic ALP interactions are related, including running effects above and below the QCD scale and provide a comprehensive analysis of the sensitivity of current and future experiments. We identify the reach of different experiments for the case of ALP dark matter and comment on how it could be distinguished from the case where it is not the dark matter. We present novel search strategies to observe quadratic ALP interactions via fifth force searches, haloscopes, helioscopes and quantum sensors. We emphasize the nonlinear behaviour of the ALP field close to the surface of the earth and point out which experimental results are independent on the local background field value.
@@ 1 Introduction
Quantum sensors have significantly enhanced their sensitivity over the past decade, creating new opportunities to explore fundamental physics questions. These advancements are particularly valuable in testing the effects of light new physics through more precise quantum sensor experiments Safronova:2017xyt ; Antypas:2022asj . A prime example of this light new physics is pseudoscalars, which emerge in theories where an approximate global symmetry is spontaneously broken Peccei:1977hh ; Peccei:1977ur ; Weinberg:1977ma ; Wilczek:1977pj ; Abbott:1982af ; Dine:1982ah . These are referred to as pseudo-Nambu Goldstone bosons, or more commonly, axion-like particles (ALP). The goal of this paper is to translate already performed and anticipated future precision measurements with quantum sensors into a sensitivity range for ALPs. We take the full effective ALP Lagrangian into account, include higher order effects when relevant and perform a consistent calculation of the axion couplings at different scales accounting for renormalisation effects. We stress the importance of quadratic ALP couplings that are formally subleading, but induce spin-independent effects that are strongly constrained from experiments that have not been considered for pseudoscalar interactions.
@@ 2.2 Quadratic ALP Interactions
where we introduced $\hat{m}=(m_{u}+m_{d})/2$ and $\Delta_{m}=(m_{u}-m_{d})/2$ to make it easier to identify the isospin-breaking terms and $\tau_{a}=m_{a}^{2}/m_{\pi}^{2}$. The sign of $\delta_{\pi}(a)$ plays a crucial role in determining the environmental effects of a massive body influencing the ALP field value. Our result (13) agrees with Kim:2022ype ; Kim:2023pvt and Beadle:2023flm in the limit $m_{a}\to 0$.
@@
$\alpha^{\text{eff}}(a)$ $=\Big{(}1+\delta_{\alpha}(a)\Big{)}\alpha\,\qquad\text{with}\qquad\delta_{\alpha}(a)=\frac{1}{12\pi}\left(1-32c_{1}\frac{m_{\pi}^{2}}{M_{N}}\right)\,\delta_{\pi}(a)\,.$ (24)
We note that Beadle:2023flm find the opposite sign in (23). Shift symmetry breaking in quadratic ALP-lepton couplings is a loop effect because the tree-level diagram in the lower row of Figure 1 results in a momentum suppressed contribution to a higher order operator with additional derivatives $(\partial a)^{2}$. The leading contribution to the quadratic ALP electron coupling is a result of photon loops Kim:2023pvt
@@ 2.5 Quadratic ALP couplings and the ALP background field
$\left(\partial_{t}^{2}-\Delta+\bar{m}_{a}^{2}(r)\right)a=J_{\text{source}}(r)+\mathcal{O}(a^{3}/f^{3})$ (45)
where $J_{\text{source}}(r)$ is suppressed by CP violating parameters or proportional to $\gamma_{5}$, and higher order terms in $a/f$ are neglected. The effective mass term can be written as
@@
$a(r,t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{a}}\cos\big{(}m_{a}t\big{)}\bigg{[}1-Z(\delta_{i})J_{\pm}\left(\sqrt{3|Z(\delta_{i})|}\right)\frac{R_{\text{source}}}{r}\bigg{]}\,,$ (47)
such that the ALP field far from the source is given by (37) (with $\beta\approx 0$), and
@@
This can be inferred from the form of the pion mass shift $\delta_{\pi}<0$ as given in (13), which is strictly negative for any sign of $c_{GG}$. In the presence of ALP-quark couplings, this is not guaranteed, but any such contribution is heavily suppressed by the small ALP mass. As a consequence, it follows from (18), (22), (24) and (26) that for all coefficients $\delta_{i}<0$, $i=\pi,M_{N},\Delta M,\alpha,e$. At the same time, the charges in (2.3) are positive in the case of an ALP interacting with gluons, for which the sole negative contributions are heavily suppressed since the ‘dilatonic’ charges $Q_{\hat{m}},Q_{\Delta M}\gg Q_{\alpha},Q_{e}$ [^3] This remains the case for the opposite sign of $\delta_{\alpha}$ found by Beadle:2023flm . Note that this is different for a scalar with quadratic couplings or even a more general model in which the sign of (51) is model-dependent. Note also, that (47) depends on the boundary conditions of the ALP field Banerjee:2022sqg .
@@
As a result, the ALP field (47) has non-zero field values close to the surface of the earth with crucial consequences for measurements of varying fundamental constants that are sensitive to the ALP field value squared, as well as for experiments measuring nuclear magnetic resonances, electric dipole oscillations Hill:2015kva ; Chu:2018avg and atomic spin precession in the ALP background field Graham:2017ivz ; Abel:2017rtm ; Agrawal:2022wjm ; Dror:2022xpi . In contrast to quadratically interacting scalars however, fundamental constants cannot change sign in this parameter space Stadnik:2021qyf ; Banerjee:2022sqg .
@@
For large ALP field values instead the full ALP potential needs to be taken into account and higher-order operators regulate the divergence in (47). Instead of (45), the full equations of motion then reads
$(\partial_{t}^{2}-\Delta+m_{a}^{2})\,a$ $=-\sin\left(\frac{a}{f}\right)\sum_{i}\,\frac{Q_{i}^{\text{source}}\delta_{i}}{f}\rho_{\text{source}}(r)$ (54)
$=\left(-\frac{a}{f}+\frac{a^{3}}{6f^{3}}-\ldots\right)\sum_{i}\,\frac{Q_{i}^{\text{source}}\delta_{i}}{f}\rho_{\text{source}}(r)\,,$ (55)
and the periodicity of the ALP potential implies a maximal field value of $a\sim\pi f$. For the case of a generic ALP field that is not a dark matter candidate the numerical solution for the full, periodic ALP potential has first been discussed in the context of new constraints from the resulting long-range forces between neutron star inspirals Hook:2017psm and a detailed analysis of the axion phase diagram in systems with finite baryonic density has been performed in Balkin:2020dsr . A crucial difference between these results and the situation considered here is the boundary condition for the ALP field far from the source. If the axion is not dark matter it vanishes for $r\to\infty$, whereas for ALP dark matter the boundary condition is given by (37). One consequence of this boundary condition is that the dependence of (53) on the ALP mass is fixed and the constraint (52) is ALP mass independent. We nevertheless expect the solution to the equation of motions of the full potential to follow the general argument that the vacuum solution (37) becomes unstable once the axion mass in the medium is tachyonic  Hook:2017psm
$m_{a}^{2}f^{2}+\frac{\sigma}{M_{N}}\rho_{N}\,\delta_{\pi}<0\,,$ (56)
[…]
$\frac{1}{r}\lesssim\Big{|}m_{a}^{2}+\frac{\sigma}{M_{N}}\rho_{N}\,\frac{\delta_{\pi}}{f^{2}}\Big{|}^{1/2}\,.$ (57)
Given these conditions, the QCD axion for which the axion mass is not a free parameter can only deviate from the vacuum solution for densities larger than the nucleon saturation density Balkin:2023xtr . If the ALP mass is treated as a free parameter the parameter space in which the ALP field is sourced by a massive body corresponds to ALP masses below the QCD axion mass, as shown in the right panel of Figure 2. An explicit example of such a model has been discussed in DiLuzio:2021pxd ; Balkin:2022qer .
Figure 2: Bounds from fifth forces on ALP-gluon coupling at the UV scale. Left panel: Constraints from searches for fifth forces induced by ALP exchange in vacuum. Right panel: Parameter space for which the ALP has non-perturbative field values.
@@ 3.1 Fifth forces
Another important effect is the acceleration experienced by test masses close to massive bodies like planets, e.g. tests of violations of the equivalence principle. Forces caused by axion fields in vacuum are captured by the potentials (59) and (62). The acceleration of a test mass $m_{T}(a)$ in the ALP background field $a(x)$ follows from the equations of motion from the low energy Lagrangian
@@
Tests of the equivalence principle are particularly sensitive to non-universal ALP couplings. We stress again that parts of this parameter space is excluded due to the nonlinearity of the ALP field close to earth which are shown for the case that the axion isn not dark matter in the right panel of Figure 2. Future tests with muonic atoms could significantly improve the sensitivity of fifth-force searches Stadnik:2022gaa .
In Figure 2 we compare constraints from different searches for fifth forces induced by ALP exchange assuming a model with one coefficient, $c_{GG}$, in the UV. The left panel shows the constraints from searches for so-called Casimir-less forces shown in blue Klimchitskaya:2015zpa ; Bauer:2023czj . The dominant contribution to this force is induced by the spin-independent potential (62). Torsion balances probe the same spin-independent force and the constraint from Adelberger:2006dh is shown in orange. Spin-dependent (dipole-dipole) interactions have been probed with comagnetometers using spin-polarised K and 3He atoms Vasilakis:2008yn and 87Rb Almasi:2018cob . The purple region labeled Ramsay scattering is excluded from measurements of spin-dependent forces in molecular H2 Ramsey:1979bzw . Future measurements of spin-dependent interactions exploiting nuclear magnetic resonances (NMR) can significantly improve the current sensitivity as shown in green dashed line in Figure 2 Arvanitaki:2014dfa ; Aybas:2021nvn . Note, that nuclear magnetic resonances are sensitive to CP-violating ALP couplings. In the case the ALP constitutes dark matter, the strongest constraint from fifth force searches are obtained by searches for violations of the equivalence principle. The MICROSCOPE collaboration measured the force necessary to keep two test masses in relative equilibrium on a satellite orbiting earth  Touboul:2017grn ; Touboul:2022yrw . The corresponding bound is shown in gray in Figure 2, but we emphasize that this parameter space is affected by the non-perturbative axion field value discussed in Section 2.5. In the right panel of Figure 2 shows these constraints from non-perturbative ALP field values following Hook:2017psm ; Balkin:2020dsr .
@@ 4.1 Quantum Clocks
where $\delta_{p}(a)$ is defined similar to $\delta_{N}(a)$ in (18) and denotes the shift in proton mass. Plugging (13), (18), (24) and (26) into (79), one can express $\delta\nu_{A/B}/\nu_{{A/B}}$ in terms of the low-energy Lagrangian parameters and the ratio $a^{2}/f^{2}$. For an oscillating ALP dark matter field, the quadratic ALP field evolves as
@@ 4.1.1 Current limits: microwave and optical clocks
Limits are obtained from the Rb/Cs atomic fountain clock, which compares the transition frequencies between different hyperfine levels in the two ground state atoms 87Rb and 133Cs. The frequency ratio measurements are sensitive to the variations in all three fundamental quantities in (79Hees:2016gop ; Flambaum:2023bnw . The measurements are also sensitive to very low frequencies corresponding to ALP mass $\sim 10^{-20}$ eV and below, due to the long time-span of the experiment. Throughout the frequency range of the experiment, the coherence time of the oscillation, $\tau_{a}=2\pi/\omega/(v/c)^{2}$ ($v/c\sim 10^{-3}$, $v$ is the virial velocity of dark matter in the galaxy) remains larger than the total timescale of measurement $T$. Although $T<<\tau_{c}$ maintains a coherent signal throughout the experiment, it also induces stochastic fluctuations in the dark matter amplitude Centers:2019dyn . Therefore, it is no longer accurate to assume the oscillation amplitude as simply $\sqrt{2\rho_{\rm DM}}/m_{a}$ because this will rather be a random variable with a sampling probability as determined by the distribution of the stochastic ALP field. This leads to corrections on the dark matter coupling that can vary from a factor of 2 up to 10.
We show the bounds from Rb/Cs in Figure 6 in the ${c_{GG}/f}$ ( UV scale) vs. $m_{a}$ plane where the coverage is for $10^{-24}\,\text{-}\,10^{-18}\ {\rm eV}$ ALP mass [^6] For dark matter below $\sim 10^{-22}$ eV, cosmological constraints prevent the assumption that the dark matter density of the Universe is entirely due to the ALP field because the de-Broglie wavelength of the ALP exceeds the size of a dwarf galaxy 2168507 . Therefore the low-frequency limits should be rescaled because $\rho_{\rm DM}$ should be a fraction of the total dark matter density. We show the limits assuming ALP constitutes 100% of dark matter density.. The constant shift in the quadratic variation in the axion field is ignored Kim:2022ype . The sensitivity coefficients $k_{\alpha},\,k_{e},\,k_{q}$ and the frequency comparison data are obtained from Ref. Hees:2016gop . The local dark matter density is assumed to be ${\rm\rho_{DM}=0.4\ GeV/cm^{3}}$ [^7] The bounds are shown without including the stochastic fluctuation factor appearing in dark matter amplitude as the exact correction factor is unclear for a pseudoscalar dark matter..
@@ 4.3 Optical/laser Interferometers
$\frac{\delta n(t)}{n}$ $\approx 10^{-2}\bigg{(}\frac{\delta\omega(t)}{\omega_{0}}-[2\delta\alpha+\delta_{e}+(\delta_{e}-\delta_{p})/2-0.024(\delta_{q}-\delta_{p})]\,\frac{\rho_{\text{DM}}}{m_{a}^{2}f^{2}}\cos\,(2\omega_{a}t)\bigg{)}$ (86)